# Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence

### v1

@@ 4.1 Supervised fine-tuning experiment
We apply uncertainty distillation to T5-base (Raffel et al., 2020) finetuned on a portion of SQUAD. Using a comparatively old and small model ensures that there will be model uncertainty on a reasonably simple dataset like SQUAD. We use defaults for most hyperparameters; in §missing 6, we discuss how uncertainty distillation is impacted by various method-specific hyperparameters.
4.2 Instruction tuning experiment
Dataset
While it is useful to be able to accurately extract confidences from a task-specific language model, in practice modern LLMs are rarely fine-tuned to any specific task. Instead, they are trained to follow instructions such that, in theory, they can perform any task that can be specified in natural language without having necessarily seen that task during training. Thus, it is important that our method can teach a language model to correctly verbalize confidences for a previously unseen task.
To evaluate this, we leverage the Super-NaturalInstructions dataset (SNI; Wang et al., 2022). We select $15$ English Q&A tasks with short-form answers. We focus on Q&A tasks for which a single correct answer exists (e.g. multiple choice problems, short-form span extraction, math problems, etc.) and thus for which correctness of a model’s prediction can reliably and efficiently be computed after normalizing lexical forms without resorting to methods such as LLM verification.
Model
We instruction-tune a T5-Large model on a remaining subset of the English tasks in the SNI dataset, making sure to explicitly hold out the $15$ tasks we use in our uncertainty distillation experiments. The result is an instruction-tuned model which we refer to as Instruct-T5, capable of performing our target Q&A tasks without having seen these tasks during training. In other words, the samples we obtain from this model do not require Instruct-T5 to be pre-trained on that specific task. See Appendix C for more details on our data selection and instruction-tuning. We train and evaluate uncertainty distillation on the combined dataset of these tasks and report the performance over the metrics described in §missing 4.3.
4.3 Metrics
We report the area under the receiver operating characteristic curve (AUROC), [^7] Calculated using scikit-learn 1.5.2 which represents the probability that a randomly chosen correct answer will be in a higher-confidence bin than a randomly chosen incorrect answer. This metric is well established in previous literature (see e.g., Hu et al. (2023)), and compares the relative rather than absolute probabilities, which allows us to use it effectively with discrete verbalized confidences. We plot the percentage of accurate answers in each bin to demonstrate that our model’s predictions correspond well with accuracy. We also report overall model accuracy, to evaluate the tradeoff between accuracy and calibration. Finally, we report accuracy of high-confidence predictions (in “very high” and “high” bins). As an established use-case for verbalized confidences is to reject lower-confidence predictions, this provides information about how useful the LLM’s predictions are. [^8] We do not report Expected Calibration Error (ECE) (Nixon et al., 2019), as it requires comparing a continuous score to the prediction’s true label, while we have five discrete bins without an obvious corresponding numerical score. Furthermore, it requires the choice of several hyperparameters which can have a large impact on performance.
4.4 Baseline
As our tasks are relatively short (with an average of fewer than five tokens in an answer), we report a lexical baseline to compare to our semantic method. [^9] In Appendix A, we also experiment with another baseline based on semantic clustering Kuhn et al. (2023). As it requires sampling 20 answers at inference time, we do not compare it here to methods that generate a single answer. Despite that, we find that it fails to replicate our performance. We extract token-level probabilities from the language model on our training and calibration splits and use this to train an isotonic regression model to calibrate the average token-level probability for each answer. [^10] We use the average probability rather than the sequence probability to normalize over different lengths, as Kuhn et al. (2023) find this improves performance. We find that using this calibrator to map the average calibrated token probability of our test set to a confidence bin (e.g. 90% to “high confidence”) is successful on our tasks, allowing us to compare our semantic method to a strong baseline.
@@ 4.4 Baseline
Table 1: AUROC and accuracy metrics for T5-base, trained on SQUAD, and Instruct-T5, trained on a combination of factual QA tasks. We find that models trained with uncertainty distillation match or outperform lexical uncertainty in verbalizing confidences on SQUAD-T5 and Instruct-T5.
5 Results and discussion
Figure 2 shows our results comparing uncertainty distillation to the lexical uncertainty baseline in terms of average accuracies in each confidence bin. In plots like this, the goal is to observe a diagonal trend line where outputs reported to have high confidence indeed have high accuracy, and those in the low confidence bins have lower accuracy. In the SQUAD experiment, we find that both uncertainty distillation and the lexical baseline fit the diagonal trend very well, showing that the methods work as intended. In the SNI experiment, the diagonal trend also holds up reasonably well, with minor exceptions at the very high/low bins.
Table 1 summarizes these plots in terms of AUROC score. Our uncertainty distillation achieves AUROC of 0.805 in the T5-base SQUAD experiment, slightly outperforming lexical baseline’s AUROC of 0.771. In the Instruct-T5 SNI experiment, uncertainty distillation achieves 0.771 AUROC compared to 0.667 from lexical baseline. We conclude that uncertainty distillation is an effective alternative to the lexical baseline, even on short-form tasks where lexical uncertainty can be expected to perform well. The advantage of uncertainty distillation in practice is that the language model is able to accurately communicate its own uncertainty without relying on separate models or output token probabilities, matching the lexical baseline without using the same mechanism or inputs. We find that the verbalized confidences produced by uncertainty distillation are highly interpretable, with high correspondence between accuracy of answers within a bin and that bins confidence.
The table also reports the accuracy of the highest confidence bin and the overall accuracy across all bins. While AUROC is the main metric for assessing calibration performance, accuracy is also useful for understanding the nuances of the result. Note that overall accuracy of T5-base on SQUAD dropped from 0.811 to 0.711 after performing uncertainty distillation; on the other hand, overall accuracy of Instruct-T5 on SNI rose from 0.387 to 0.449. Some changes in model accuracy are expected due to the fine-tuning step described in §missing 3.3; we discuss the potential tradeoffs between accuracy and calibration further in §missing 7.
In more detail, for the SNI QA tasks we find that uncertainty distillation results in underconfidence in the lowest bin– the accuracy in the “very low” bin is higher than the ideal range for that bin (between 0 and 20% accuracy). This is likely due to the additional fine-tuning performed in uncertainty distillation. As Instruct-T5 has not been trained on the tasks before sampling, additional fine-tuning improves its performance on these tasks (as can be seen the increased overall accuracy). The model is more likely to get an answer right after fine-tuning, but the annotations do not reflect that change, resulting in underconfidence in those bins. Conversely, the lexical baseline is overconfident in the highest bin, with the accuracy falling well under the ideal range (between 80% and 100% accuracy). This may indicate that token probabilities are a less reliable measure of uncertainty when the language model has not been previously trained on in-distribution training data.
Figure 2: Average accuracy within each confidence bin for our main experiments. We find that our confidence bins are relatively reasonable, with low-confidence bins containing a low proportion of correct answers and high-confidence bins containing a high proportion.
6 Analysis
Model Method AUROC Overall Accuracy High Accuracy
FLAN-T5 Uncertainty distillation 0.873 0.614 0.875
Lexical baseline 0.892 0.657 0.912
Table 2: AUROC and accuracy metrics when using FLAN-T5, which does not have an unseen calibration set. We find that our semantic uncertainty method performs equivalently to a lexical uncertainty method on FLAN-T5.
6.1 Assumption of calibration set
In §missing 3, we assume that we have access to held-out calibration data. However, due to the opacity and size of pretraining datasets, it is increasingly challenging to guarantee that this assumption holds. We test our method on FLAN-T5 Chung et al. (2022), an instruction-tuned model trained on a dataset containing our tasks. Importantly, we not only verify that Flan-T5 has been instruction-tuned on our tasks, but has seen samples from the calibration set of our test tasks. This allows us to investigate the effect of data contamination on calibration of verbalized confidences.
We report our results in Table 2. We find that while uncertainty distillation still produces meaningful confidence bins, it no longer outperforms lexical uncertainty. We conclude that our method works in the absence of this assumption, but not as effectively as token-level probabilities, which are likely well-calibrated due to the model’s previous training on these examples. We discuss results for FLAN-T5 further in §missing 6.3 and §missing 6.4, and find that its behavior differs significantly from results on models where we have an unseen calibration set.
6.2 Number of samples
@@
Our Monte Carlo estimation of probability requires sampling repeatedly from a model before normalizing and calculating probability. In Figure 3, we show that the number of samples used to estimate the initial probabilities has a significant impact if chosen to be too low; however, there are diminishing returns as the number of samples increases. We choose to use 1,000 samples in all of our experiments, as more than that is unlikely to achieve anything but marginal improvement.
@@
6.3 Ablation of post-hoc calibration
If the model’s initial predictions are poorly calibrated, the post-hoc calibration step should help to better align probabilities in the training data with the true likelihood of success. In Table 3, we show the results of models trained with and without this post-hoc calibration step. We find no apparent benefit of post-hoc calibration for T5-base or Instruct-T5. These models are already well-calibrated on their respective domains (see Appendix B), and thus the post-hoc calibration step does not have a significant impact; had the model probabilities been worse-calibrated, the post-hoc calibration step would have increased importance. In the case of FLAN-T5, post-hoc calibration decreases AUROC. Notably, FLAN-T5 does not have unseen calibration data, while Instruct-T5 and T5-base are both trained excluding calibration data. This suggests that in cases when unseen calibration data cannot be obtained, uncertainty distillation may be more effective without the post-hoc calibration step.
Model With Post-hoc No Post-hoc
T5-base 0.804 0.800
Instruct-T5 0.751 0.751
@@ 6.3 Ablation of post-hoc calibration
Table 3: AUROC of models with and without post-hoc calibration at training time. We find that there is no notable performance increase on these datasets with post-hoc calibration.
6.4 Adding incorrect examples
While adding incorrect examples into the training data has the potential to provide more examples at different levels of confidences, it also is likely to increase the likelihood that a model generates an incorrect answer. To demonstrate this effect, in Table 4, we show the AUROC and accuracy for models trained with different amounts of incorrect samples. For T5-base and Instruct-T5, we find that adding only two incorrect samples per correct sample dramatically increases AUROC while decreasing accuracy. While this would seem to indicate a fundamental tradeoff between accuracy and calibration, we find that the same is not as obviously true for FLAN-T5; while the accuracy may decrease and AUROC may increase, the effects are not as significant as they are for T5-base or Instruct-T5. One possible interpretation of this is that its predictions are shaped by the fact that the data was included in its instruction-tuning corpus, leading to less dramatic shifts when trained.
While adding incorrect samples may improve AUROC, it the number of training examples by a factor of the number of incorrect examples added (e.g. a training set with 100 examples would train on 100 augmented answers with 0 incorrect examples added, 200 augmented answers with one incorrect example added, etc.) This leads to increased compute at training time. For this reason, in addition to the decreased accuracy, we recommend adding a low number of incorrect examples to the training dataset, and in our experiments limit to two incorrect answers per question.
0 1 2 3
T5-Base
AUROC 0.707 - 0.805 -
Accuracy 0.791 - 0.711 -
Instruct-T5
AUROC 0.723 - 0.751 -
Accuracy 0.529 - 0.449 -
FLAN-T5
AUROC 0.868 0.876 0.873 0.883
Accuracy 0.609 0.620 0.614 0.611
Table 4: AUROC of models trained with varying numbers of incorrect examples allowed per question. There is a general trend towards increasing AUROC and decreasing accuracy when incorrect examples are included, although this is less pronounced for FLAN-T5. FLAN-T5 displays only a slight version of this trend; we evaluate over more points to demonstrate that, while slight, the same phenomenon occurs.
7 Conclusion
Findings
We find considerable evidence that using uncertainty distillation to verbalize semantic uncertainty can perform comparably to or better than lexical uncertainty in QA tasks, despite operating using an entirely different mechanism. We demonstrate that our method is successful in all three settings we test it in– a small fine-tuned model, a large instruction-tuned model with unseen data, and an instruction-tuned model calibrated with information in its instruction-tuning set. We also find that in cases where the initial model has not been tuned on the specific task before uncertainty quantification methods are used, the lexical uncertainty baseline tends towards overconfidence, while uncertainty distillation tends to be underconfident.
One high-level takeaway is that there appears to be a tradeoff between an LLM’s ability to predict its own confidence and overall model accuracy. In our experiments, interventions that improve AUROC decrease accuracy and vice versa. While some amount of this may be remedied by increased hyperparameter tuning or small operations, it is also plausible that there are some fundamental tradeoffs.
Future work
In this work, we demonstrate that uncertainty distillation exhibits strong performance comparable to a lexical baseline only on tasks where the lexical method succeeds. Future work should examine its effectiveness in cases where a lexical baseline fails. We focus on strings that are quickly and easily normalized with preprocessing functions, but our method could easily be applied to tasks outside simple QA through the use of LLM verifiers to calculate binary correctness. Future work may also investigate the robustness of the model’s internal representation of uncertainty to domain shifts, particularly in the case of multi-task generalization—do the confidence scores remain interpretable and well-calibrated on similar unseen tasks without further training?
Limitations
While we anticipate that increasing model size would only increase our method’s efficacy, we do not test on models larger than T5-large, both for computational reasons and because we cannot ascertain which data is in training sets for many larger models. We also run on relatively simple QA tasks, where computing uncertainty is comparatively straightforward. Our method is also not infallible, with imperfect accuracies in high-confidence bins– incorrect high confidence predictions may present a small risk of prompting users to trust predictions more than they otherwise would have.
Appendix A Comparison to semantic clustering
In Table 5, we compare uncertainty distillation to semantic clustering as described by Kuhn et al. (2023). In this method, twenty answers are sampled at inference time, and DeBERTaHe et al. (2020) is used to determine how many are semantically equivalent; the most common answer is evaluated, and the proportion of answers in that cluster is considered to be the probability. Similarly to our own method and the baseline, we evaluate AUROC on the binned scores. We find that semantic clustering fails to outperform uncertainty distillation despite using 20 times the compute, and additionally find that overall accuracy with semantic clustering is highly variable, though accuracy in the top bins remains high.
Model Method AUROC Overall Accuracy High Accuracy
Instruct-T5 Uncertainty distillation 0.751 0.449 0.839
Semantic clustering 0.708 0.65 0.917
FLAN-T5 Uncertainty distillation 0.873 0.614 0.875
Semantic clustering 0.868 0.346 0.879
Table 5: AUROC and accuracy metrics compared to semantic clustering. We find that uncertainty distillation compares favorably with a single sample, rather than the 20 samples required at inference time by semantic clustering.
Appendix B Model calibration on QA tasks
While LLM overconfidence is well-documented on many tasks, we find that our LLMs are well-calibrated on the short QA tasks we use for uncertainty estimation. Figure 4 shows the reliability diagrams for T5-base on SQUAD and Instruct-T5 on SNI. The models’ predicted confidences align well with their actual accuracies; a post-hoc calibrator will not significantly alter the output probabilities.
@@ Appendix B Model calibration on QA tasks
Appendix C SuperNatural-Instructions Tasks
C.1 Target Calibration Tasks
@@
C.2 Instruction-Tuning Tasks
Because most modern instruction-tuned models are trained on all of Super-NaturalInstructions, they have seen the our calibration target tasks during instruction-tuning. Therefore, we instruction-tune our own T5 model to test the effectiveness of our method on unseen tasks. Our model is trained on a subset of the SuperNatural-Instructions dataset (Wang et al., 2022). Specifically, we instruction-tune on the English split used in the original paper but we take out our target calibration tasks identified in §missing C.1. This gives us a training dataset of $879$ instruction-tuning tasks, with a total of roughly 1.2M training samples total.
@@ C.2 Instruction-Tuning Tasks
Appendix D Instruction-Tuning T5
@@
Appendix E Resource Reporting
E.1 Compute Resources
Here we report the compute resources used in this work. Instruction-tuning T5 took a total of $200$ GPU hours across $4$ NVIDIA-V100s. Running uncertainty distillation on Instruct-T5 and FLAN-T5 took 16 hours per model on a single NVIDIA-H100. Finetuning T5-base on SQUAD for our initial model took 3 hours on a single NVIDIA RTX 2080, and training using uncertainty distillation took 8 hours on a single NVIDIA-V100. Our lexical baseline for SQUAD took one hour on one NVIDIA RTX 2080, and for SNI took three hours on one NVIDIA RTX 2080. Sampling for SQUAD took a total of 60 GPU hours on NVIDIA-V100s, and for SNI took 45 GPU hours on NVIDIA-A100s.
E.2 Resource Intended Use
@@
Appendix F Uncertainty distillation hyperparameters
In Table 7, we show the training hyperparameters for uncertainty distillation training.
@@ Appendix F Uncertainty distillation hyperparameters
Table 7: Hyperparameters for training all models but Instruct-T5 (see Appendix D for details). All models are trained with the AdamW optimizer.
Appendix G Algorithm

### v2

@@ Appendix A Uncertainty distillation on supervised fine-tuned models
We apply uncertainty distillation to T5-base (Raffel et al., 2020) finetuned on a portion of SQUAD. We use defaults for most hyperparameters, and report hyperparameters in Appendix I.
Results
Table 4 shows the results on the fine-tuned T5-base model. Uncertainty distillation achieves AUROC of 0.805 in the T5-base SQUAD experiment, slightly outperforming the lexical baseline’s AUROC of 0.771.
@@ Results
Table 4: AUROC and accuracy metrics for T5-base, trained on SQUAD. We find that even in this setting, a model trained with uncertainty distillation outperforms lexical uncertainty in verbalizing confidences on SQUAD-T5
Appendix B Number of samples
@@
Our Monte Carlo estimation of probability requires sampling repeatedly from a model before normalizing and calculating probability. In Figure 3, we show that the number of samples used to estimate the initial probabilities has a significant impact if chosen to be too low; however, there are diminishing returns as the number of samples increases. We therefore choose to use 1,000 samples in all of our experiments with FLAN-T5 and Instruct-T5, as more than that is unlikely to achieve anything but marginal improvement.
@@
Appendix C Prompts
C.1 Mistral, Llama
Prompt baselines, uncertainty distillation
Answer the following question and state confidence in the answer (very low, low, medium, high, very high). Enclose concise reasoning in <reasoning> </reasoning> tags, confidence in <confidence> </confidence> tags, and your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
A) 24
B) 35
C) 42
D) 50
<reasoning> This can be solved with multiplication. The answer is 7*6, or 42.</reasoning> <answer> C) 42 </answer> <confidence>very high</confidence>." Your answer should not include words.
Sampling, lexical baseline
Answer the following question. Enclose concise reasoning in <reasoning> </reasoning> tags and your FINAL answer in <answer> </answer> tags without any of your work, like this: "If each of Lisa’s 7 chickens lays 6 eggs, how many eggs does Lisa have?
A) 24
B) 35
C) 42
D) 50
<reasoning> This can be solved with multiplication. The answer is 7*6, or 42.</reasoning> <answer> C) 42 </answer>." Your answer should not include words.
C.2 Instruct-T5, FLAN-T5
Each task in SNI has an associated instruction. For sampling and the lexical baseline, we simply use this instruction. For uncertainty distillation, we append ‘‘Additionally state how confident you are in your answer’’ to the instruction.
Appendix D Effects of post-hoc calibration on well-calibrated models
If the model’s initial predictions are poorly calibrated, the post-hoc calibration step should help to better align probabilities in the training data with the true likelihood of success; indeed, we find that post-hoc calibration has a significant positive effect with our largest models. For instance, Llama-3B on SocialIQA achieves 0.784 AUROC when trained on post-hoc calibrated data, and only 0.680 when identically trained on data without post-hoc calibration.
However, how does post-hoc calibration impact the model when the model is already well-calibrated on the specific task, or when the model has previously been trained on the calibration data? Figure 4 shows the reliability diagrams for T5-base on SQUAD and Instruct-T5 on SNI. The models’ predicted confidences align well with their actual accuracies; this allows us to investigate whether post-hoc calibration has a significant impact on AUROC. Additionally, FLAN-T5 has been previously tuned on our calibration set; this gives us a setting to investigate the impact of post-hoc calibration when unseen calibration data is unavailable.
In Table 5, we show the results of the smaller models trained with and without this post-hoc calibration step. We find no apparent benefit of post-hoc calibration for Instruct-T5 or fine-tuned T5-base. These models are already well-calibrated on their domains; therefore, a post-hoc calibrator does not significantly alter the output probabilities.
In the case of FLAN-T5, post-hoc calibration decreases AUROC. This suggests that in cases when unseen calibration data cannot be obtained, uncertainty distillation may be more effective without the post-hoc calibration step.
Dataset Model With Post-hoc No Post-hoc
SQUAD T5-base 0.804 0.800
SNI Instruct-T5 0.751 0.751
@@ Appendix D Effects of post-hoc calibration on well-calibrated models
Table 5: AUROC of well-calibrated models with and without post-hoc calibration at training time. We find that there is no notable performance increase with post-hoc calibration, and that there is a performance decrease when the model has previously been tuned on the calibration data.
@@
Appendix E SuperNatural-Instructions Tasks
E.1 Target Calibration Tasks
@@
E.2 Instruction-Tuning Tasks
Because most modern instruction-tuned models are trained on all of Super-NaturalInstructions, they have seen the our calibration target tasks during instruction-tuning. Therefore, we instruction-tune our own T5 model to test the effectiveness of our method on unseen tasks. Our model is trained on a subset of the SuperNatural-Instructions dataset (Wang et al., 2022). Specifically, we instruction-tune on the English split used in the original paper but we take out our target calibration tasks identified in §missing E.1. This gives us a training dataset of $879$ instruction-tuning tasks, with a total of roughly 1.2M training samples total.
@@ E.2 Instruction-Tuning Tasks
Appendix F Ministral plots
In Figure 5 we display the plots with Ministral-8B. As reflected in the AUROC score in Table 1, calibration is slightly worse; however, compared to baselines, it still does a more accurate job of forecasting confidence.
Figure 5: Average accuracy within each confidence bin for our main experiments. We do not plot bins with fewer than 10 samples.
Appendix G Instruction-Tuning T5
@@
Appendix H Resource Reporting
H.1 Compute Resources
Here we report the compute resources used in this work. Instruction-tuning T5 took a total of $200$ GPU hours across $4$ NVIDIA-V100s. Running uncertainty distillation on Instruct-T5 and FLAN-T5 took 16 hours per model on a single NVIDIA-H100. Finetuning T5-base on SQUAD for our initial model took 3 hours on a single NVIDIA RTX 2080, and training using uncertainty distillation took 8 hours on a single NVIDIA-V100. Finetuning Ministral-8B (LoRA) and finetuning Llama-3B each took took three hours on two NVIDIA-A100s. Our lexical baseline for SQUAD took one hour on one NVIDIA RTX 2080; for SNI took three hours on one NVIDIA RTX 2080; for MMLU took three hours on one NVIDIA-A100; and for SocialIQA took thre hours on one NVIDIA-A100. Prompting for MMLU and prompting for SocialIQA took 1 hour on one NVIDIA-A100. Sampling for SQUAD took a total of 60 GPU hours on NVIDIA-V100s; for SNI took 45 GPU hours on NVIDIA-A100s; for SocialIQA took 350 hours on NVIDIA-A100s; and for MMLU took 350 hours on NVIDIA-A100s.
H.2 Resource Intended Use
@@
SocialIQA Sap et al. (2019) is not explicitly licensed, but they state that they “ establish Social IQa as a resource” for future models.
MMLU Hendrycks et al. (2020a) is published under the MIT license, which allows users to freely copy, use, and change the licensed material.
Appendix I Uncertainty distillation hyperparameters
In Table 7 and Table 8, we show the training hyperparameters for uncertainty distillation training. All experiments in Table 7 added two incorrect answers per question, and in Table 8 added one incorrect answer per question.
@@ Appendix I Uncertainty distillation hyperparameters
Table 7: Hyperparameters for training all T5 models but Instruct-T5 (see Appendix G for details). All models are trained with the AdamW optimizer.
Model Epochs Learning rate Batch size LoRA rank LoRA alpha
Llama-3B/MMLU 3 4e-5 4 - -
Llama-3B/SocialIQA 1 3e-5 4 - -
Ministral-8B/MMLU 3 5e-5 4 16 32
Ministral-8B/SocialIQA 1 3e-5 4 8 16
Table 8: Hyperparameters for training all Llama and Ministral models. Gradient accumulation steps is 1 for each model. All models are trained with the AdamW optimizer.
Appendix J Algorithm